### v2

# LaV-CoT: Language-Aware Visual CoT with Multi-Aspect Reward Optimization for Real-World Multilingual VQA

Abstract.
[…]
Vision-Language Models; Multilingual Multimodal Reasoning; Visual Chain-of-Thought; Reinforcement Tuning
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Companion Proceedings of the ACM Web Conference 2026; April 13–17, 2026; Dubai, UAEccs: Computing methodologies Machine learningccs: Computing methodologies Computer visionccs: Computing methodologies Natural language processing**footnotetext: This work is a preprint.
1. Introduction
Large Vision Language Models (VLMs) (Li et al., 2022; Alayrac et al., 2022; Liu et al., 2023; Lu et al., 2024; Wang et al., 2024a, b; Li et al., 2024b; Sun et al., 2024; Bai et al., 2025; Zhu et al., 2025; Wang et al., 2025) have advanced rapidly in recent years, demonstrating impressive performance on multimodal tasks such as image captioning (Wang et al., 2023), visual question answering (VQA) (Chen et al., 2022b), and complex reasoning (Li et al., 2024a). By integrating visual and textual modalities, these models learn rich semantic representations and enable a wide range of real-world applications. With the growing demand for global accessibility, multilingual visual question answering (mVQA) (Changpinyo et al., 2023) has emerged as a critical capability for deploying VLMs at scale.
Despite recent progress in vision language models (VLMs), achieving reliable reasoning in multilingual multimodal settings remains challenging. Textual Chain-of-Thought (CoT) (Wei et al., 2022) reasoning improves interpretability and supports step-by-step inference (Wei et al., 2022). However, when applied to multilingual visual question answering (mVQA), a text-only rationale often under-exploits visual cues and weakens visual grounding. Orthogonal to the choice of CoT, current VLMs for mVQA still exhibit persistent issues: (i) language inconsistency, where outputs fluctuate across languages; (ii) visual–textual misalignment, reflecting insufficient grounding between vision-translated text and visual content; and (iii) limited multilingual visual reasoning, especially for tasks requiring complex cross-lingual, multimodal inference (Alam et al., 2024; Maaz et al., 2024; Thawakar et al., 2025; Wu et al., 2025). Such limitations constrain the deployment of VLMs in international products, which demand both robustness and interpretable reasoning.
To address these challenges, we propose LaV-CoT, a Language-aware Visual CoT reasoning framework with Multi-Aspect Reward Optimization. As illustrated in Figure 1, LaV-CoT introduces an interpretable multi-stage reasoning pipeline that integrates four key components: (1) Text Summary with Bounding Box (BBox) (Yang et al., 2022), (2) Language Identification, (3) Spatial Object-level Captioning (Huang, 2024), and (4) Step-by-step Logical Reasoning (Surís et al., 2023; Zhang et al., 2024; Xu et al., 2025). This structured pipeline explicitly disentangles language and visual reasoning, enabling fine-grained cross-modal alignment and improving interpretability in multilingual settings.
Figure 1. Overview of Lav-CoT: (a) Direct model answers may be incorrect and exhibit language inconsistency. (b) Incorporating CoT reasoning enhances reasoning transparency but still exhibits linguistic inconsistency. (c) Our method introduces a multi-stage reasoning pipeline, yielding accurate and consistent final answers.:
A further challenge lies in constructing high-quality multilingual reasoning data. Manual annotation is costly and often infeasible at scale. To this end, we design an automatic data curation method that generates multilingual CoT annotations via iterative generation, correction, and refinement (Zelikman et al., 2022; Lightman et al., 2023; Shao et al., 2025). This scalable process produces structured and verifiable reasoning traces, ensuring that training data captures both linguistic fidelity and multimodal reasoning quality.
On the training side, LaV-CoT adopts a two-stage paradigm combining Supervised Fine-Tuning (SFT) (Li et al., 2022; Chen et al., 2022b; Bai et al., 2025) with Language-aware Group Relative Policy Optimization (GRPO) (Lu et al., 2023; Shao et al., 2024; Liu et al., 2025). Unlike standard optimization methods, our GRPO variant is guided by verifiable multi-aspect rewards—including language consistency reward, text segments and object count reward, final answer edit distance reward and format reward—yielding stable optimization and robust reasoning across languages and modalities.
[…]
Our main contributions are as follows:
  • We propose LaV-CoT, the first framework to unify language-aware visual CoT reasoning with multi-aspect reward optimization for multilingual multimodal tasks.
  • We design an automatic data curation method that produces scalable, high-quality multilingual CoT annotations through iterative generation, correction, and refinement.
  • We develop a language-aware GRPO algorithm with verifiable multi-aspect rewards, enhancing reasoning robustness and cross-modal alignment.
  • We validate our framework through extensive experiments on public benchmarks and further reinforce its effectiveness via a online A/B test, demonstrating state-of-the-art multilingual multimodal reasoning performance.
Figure 2. The framework includes an automated data generation pipeline, which leverages a multi-step reasoning process comprising (a) Text Summary with BBox, (b) Language Identification, (c) Spatial Object-level Image Captioning, and (d) Step-by-step Logical Reasoning. This reasoning pipeline first produces vanilla CoT annotations, which are then iteratively refined through rigorous verification to ensure high-quality supervision. The generated data subsequently supports a two-stage training paradigm combining SFT with GRPO. Reward computation considers language consistency, count accuracy, and the edit distance between predicted and ground-truth answers, collectively enhancing structural understanding and reinforcing robust reasoning capabilities.:
2. RELATED WORKS
2.1. Large Multilingual Vision-Language Reasoning Model
Large multilingual vision-language reasoning models bridge complex visual understanding and cross-lingual semantic reasoning. Early works focused on combining visual features with symbolic reasoning to address structured visual tasks (Andreas et al., 2016; Liang et al., 2018). With the advent of large language models (LLMs), recent approaches leverage their powerful reasoning abilities to interpret and generate multimodal content across multiple languages(Alayrac et al., 2022; Chen et al., 2022b). To address multilingual challenges, models now incorporate text encoders and alignment strategies, enabling seamless reasoning across diverse languages (Karoui et al., 2023; Zhang et al., 2023). Visual representation advances, such as cognition-focused visual tokens, boost fine-grained reasoning (Bigverdi et al., 2025; Chen et al., 2025). Innovative frameworks treat the language model as a reasoning agent interacting dynamically with specialized visual modules, supporting flexible and interpretable workflows, as in VISPROG (Asai et al., 2022). And (Ahmat et al., 2025) propose a Multi-grained Multilingual Vision-Language Pre-training (M2-VLP) model, which uses a unified multi-grained contrastive learning paradigm to integrate cross-language and cross-modal alignment.(Ranaldi et al., 2025) proposed the R2-MultiOmnia method, which guides key elements in the model’s abstract reasoning process and optimizes the reasoning trajectory through self-correction, enabling Multimodal Large Language Models (MLLMs) to maintain consistent reasoning capabilities across different languages and improve modal performance.
2.2. Chain-of-thought in Large Language Models
Chain-of-thought (CoT) prompting (Wei et al., 2022) provides a systematic, stepwise reasoning framework that helps large language models tackle challenging problems such as commonsense reasoning (Wei et al., 2022; Kumar et al., 2022) and logical deduction (Cobbe et al., 2021; Wei et al., 2022). By breaking down complex questions into a sequence of intermediate reasoning steps, CoT guides the model to produce detailed, incremental solutions rather than direct answers. Recent research has shown that incorporating CoT strategies significantly boosts the reasoning abilities of vision-language models (VLMs). For example, Prism (Chen et al., 2022a) separates the reasoning pipeline into distinct perception and inference stages to enhance understanding. MSG (Luo et al., 2023) introduces forced Chain-of-Thoughts, laying the groundwork for structured prompting methods. Approaches like Distilling CoT (Hsieh et al., 2023) and Visual Program Distillation (Hu et al., 2024) focus on transferring CoT reasoning patterns into VLMs. Visual CoT (Shao et al., 2025) enhances interpretability by producing bounding boxes that highlight relevant image regions alongside textual answers, providing clearer visual grounding of the reasoning process. Meanwhile, LLaVA-CoT (Lin et al., 2023) employs a structured Chain-of-Thought mechanism, which systematically decomposes complex reasoning tasks and thereby achieves improved accuracy in multimodal reasoning. Building upon these advancements, our work proposes a systematic and interpretable CoT framework explicitly designed to address the challenges of multilingual vision-language reasoning.
2.3. Reinformence Tuning
Reinforcement Learning (RL) (Sutton and Barto, 2018; Mnih et al., 2015; Silver et al., 2016) is a fundamental paradigm in machine learning, where an agent learns to interact with an environment by performing actions, receiving feedback in the form of rewards, and iteratively updating its policy to maximize cumulative long-term returns. Classical RL algorithms, such as Q-learning (Watkins and Dayan, 1992), have been successfully applied across diverse domains including robotics, game playing (e.g., AlphaGo), and autonomous control. With the advent of large language models (LLMs), Reinforcement Learning with Human Feedback (RLHF) (Christiano et al., 2017; Stiennon et al., 2020) has become a prominent technique to fine-tune models by leveraging human preference signals. RLHF commonly employs policy optimization algorithms like Proximal Policy Optimization (PPO) (Ouyang et al., 2022) and Direct Preference Optimization (DPO) (Rafailov et al., 2023) to steer model behavior, thereby enhancing alignment, coherence, and helpfulness in generated responses. More recently, advanced policy optimization variants such as Group Relative Policy Optimization (GRPO) (Lu et al., 2023) have been proposed, which modify the standard value-function estimator with a group-relative advantage function computed over multiple candidate responses per prompt. Moreover, RL-based fine-tuning strategies have also been successfully extended to vision-language models (VLMs). V-DPO (Xie et al., 2024) uses vision-guided Direct Preference Optimization to enhance visual context learning at training time. In our work, we design verifiable multi-aspect reward functions to enhance the granularity and reliability of feedback signals during reinforcement learning.
3. Method
Our framework enables a progressive, step-by-step reasoning process that strengthens the reasoning capabilities of Vision-Language Models (VLMs). Leveraging structured multi-stage thinking, it establishes a systematic and efficient reasoning pipeline. To ensure high-quality training data, the data curation stage incorporates iterative CoT evaluations to identify and correct errors until each reasoning sequence is fully validated. Finally, a two-stage training strategy with a novel, multi-aspect reward design is employed to obtain the final optimized model.
3.1. Multi-stage Reasoning
The multi-stage reasoning design is motivated by the way humans naturally approach image understanding: when examining a document image, we first locate salient text regions and retain their summarized content, then identify the language, recognize objects and their spatial relationships, and finally integrate all information to reason step by step toward an answer. As illustrated in Figure 2, our model performs reasoning through four structured stages:
  • (a)
    Text Summary with Bounding Boxes. For text-centric images, we first detect text segments within the image. Since extracting the complete OCR content may be costly in terms of token usage, we apply a summarization strategy to generate concise yet informative representations of the detected text. The output is formatted as a list of bounding boxes paired with summarized text, where the length of the list is later used as a reward signal during training.
  • (b)
[…]
  • (c)
[…]
  • (d)
    Step-by-Step Logical Reasoning. Leveraging the outputs from all previous steps as evidence, we first understand the given question, then devise a detailed solution plan, and finally execute the plan step-by-step until arriving at the final answer.
[…]
3.2. Dataset Curation
Most existing VQA datasets lack the detailed reasoning processes necessary to effectively train the multilingual reasoning vlm model. To address this limitation, we compile a new dataset by integrating samples from several widely used VQA benchmarks, resulting in a total of 148k image–question–CoT–answer pairs, as illustrated in Table 3.
Specifically, as shown in Algorithm 1, we start from the original question–answer triplets $\langle I,Q,A\rangle$, where $I$ denotes the image, $Q$ the question, and $A$ the corresponding answer. We first prompt a GPT-based generator $f_{\text{gen}}$ to produce an initial sequence of vanilla Chain-of-Thought (CoT) steps $T_{\text{init}}=\{s_{i}\}_{i=1}^{|T_{\text{init}}|}$. Next, we prompt an evaluator $f_{\text{eval}}$ to score each step $s_{i}\in T_{\text{init}}$. For any step whose score falls below the threshold $\tau$, we iteratively perform the following procedure: First, we apply the evaluator to the step and then locate the erroneous part, denoted as $s_{\text{error}}$, using the function $\text{Locate}(f_{\text{eval}}(s_{i}))$. Next, a corrected step $s_{i}$ is generated by applying the function Correct to $s_{\text{error}}$, i.e., $s_{i}$ is updated as $\text{Correct}(f_{\text{gen}}(s_{\text{error}}))$. The corrected step then replaces the original step in the sequence $T_{\text{init}}$. Finally, the updated step is re-evaluated to obtain the new score using $f_{\text{eval}}(s_{i})$. This evaluation–correction–update loop repeats until all steps in $T_{\text{init}}$ exceed the threshold. The final verified Chain-of-Thought sequence is denoted as $T$, and the output dataset consists of $\langle I,Q,T,A\rangle$. Detailed prompt design are shown in Appendix B.
Our dataset covers 13 languages, including English (EN), Chinese (ZH), Portuguese (PT), Arabic (AR), Turkish (TR), Russian (RU), German (DE), French (FR), Italian (IT), Japanese (JA), Korean (KO), Thai (TH), and Vietnamese (VI). These languages represent a diverse range of linguistic families. Our datasets are selected from diverse sources, including COCO2017(Lin et al., 2014), a large-scale dataset for object detection, segmentation, and image captioning containing everyday scenes with common objects in context; Visual Genome(Krishna et al., 2016), which provides densely annotated images linking objects, attributes, relationships, and region descriptions for visual reasoning; GQA(Hudson and Manning, 2019), emphasizing compositional reasoning over real-world images with complex question-answer pairs; OCR-VQA(Mishra et al., 2019) and TextVQA(Singh et al., 2019), both focusing on reading and reasoning over text within images for text-centric visual question answering; Llava-Pretrain(Lin et al., 2023), a multimodal pretraining dataset supporting various VQA and reasoning tasks across multiple domains; and MTVQA(Tang et al., 2024), a multilingual text-based VQA dataset designed to evaluate cross-lingual multimodal understanding.
Algorithm 1 Verified CoT Data Generation for Input $\langle I,Q,A\rangle$
0: Image II, Question $Q$, Answer AA, Generator $f_{\text{gen}}$, Evaluator $f_{\text{eval}}$, Threshold $\tau$
0: Verified CoT data $\langle I,Q,T,A\rangle$
1: Initialize $T_{\text{init}}\leftarrow f_{\text{gen}}(\langle I,Q,A\rangle)$
2:for $i=1$ to $|T_{\text{init}}|$ do
3:  $s_{i}\leftarrow T_{\text{init}}[i]$
4:  $score\leftarrow f_{\text{eval}}(s_{i})$
5:  while $score<\tau$ do
6:   $s_{\text{error}}\leftarrow\text{Locate}(f_{\text{eval}}(s_{i}))$
7:   $s_{i}\leftarrow\text{Correct}(f_{\text{gen}}(s_{\text{error}}))$
8:   Update $T_{\text{init}}[i]\leftarrow s_{i}$
9:   $score\leftarrow f_{\text{eval}}(s_{i})$
10:  end while
11:end for
12:return $\langle I,Q,T,A\rangle$
3.3. Model Training
We adopt Supervised Fine-Tuning (SFT) to equip the model with robust multilingual and multimodal reasoning abilities. To further refine output quality, we introduce a composite reward function, $R_{\mathrm{Multi\_Aspect}}$, which aggregates multiple rule-based criteria into a single supervisory signal. Instead of relying on an explicit value function or critic, candidate outputs are comparatively evaluated within the same group, and their raw rewards are transformed into relative advantage scores. This pairwise normalization stabilizes optimization and promotes the preference for higher-quality responses.
The multi-aspect reward $R_{\mathrm{Multi\_Aspect}}$ consists of four complementary components: three novel rewards designed to capture different aspects of multilingual visual-textual reasoning, and a default format reward that ensures the generated output adheres to the expected structural conventions.
1. Language Consistency Reward ($R_{\mathrm{Lang}}$). To encourage the model to perform reasoning in the target language, we compare the language predicted by the model with the labeled primary language of the input. Let $L$ denote the ground-truth language label and $\hat{L}$ represent the language identified by the model. The reward is defined as:
$R_{\mathrm{Lang}}=\begin{cases}1,&\text{if }L=\hat{L}\\ 0,&\text{otherwise}\end{cases}$
2. Text Segments and Object Count Reward ($R_{\mathrm{Count}}$). To ensure accurate text segmentation and object count, let $N_{Ts}$ and $\hat{N}_{Ts}$ be the numbers of reference and predicted text segments, and $N_{Obj}$ and $\hat{N}_{Obj}$ be the numbers of reference and predicted main objects. The reward is defined as:
$R_{\mathrm{Count}}=1-\frac{|N_{Ts}-\hat{N}_{Ts}+N_{Obj}-\hat{N}_{Obj}|}{N_{Ts}+N_{Obj}}$
3. Edit Distance of Final Answer Reward ($R_{\mathrm{Answer}}$). We compute the normalized Levenshtein distance D(Y,Y^)D($\hat{Y}$) between the reference answer $Y$ and the model prediction $\hat{Y}$:
$R_{\mathrm{Answer}}=1-\frac{D(Y,\hat{Y})}{\max(\ell_{Y},\ell_{\hat{Y}})}$
4. Format Reward ($R_{\mathrm{Format}}$). This default reward encourages the model to generate outputs that strictly follow the expected structural format. Let $\hat{Y}$ denote the model output, and let $T$ represent the set of required tags, such as <think></think> and <answer></answer>. The reward is defined as:
$R_{\mathrm{Format}}=\begin{cases}1,&\text{if }T\subseteq\hat{Y}\\ 0,&\text{otherwise}\end{cases}$
The final multi-aspect reward is a weighted combination of these four components:
$R_{\mathrm{Multi\_Aspect}}=\alpha\,R_{\mathrm{Lang}}+\beta\,R_{\mathrm{Count}}+\gamma\,R_{\mathrm{Answer}}+\delta\,R_{\mathrm{Format}},$
where $\alpha$, $\beta$, $\gamma$, and $\delta$ are non-negative coefficients that control the relative importance of each reward component.
The objective function of GRPO(Lu et al., 2023) is defined as:
$\mathcal{J}_{\mathrm{GRPO}}(\theta)$ $=\frac{1}{G}\sum_{i=1}^{G}\Bigg[\min\Big(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid q)}A_{i},$
$\qquad\quad\mathrm{clip}\Big(\frac{\pi_{\theta}(o_{i}\mid q)}{\pi_{\theta_{\mathrm{old}}}(o_{i}\mid q)},1-\epsilon,1+\epsilon\Big)A_{i}\Big)\Bigg]$
(1) $\quad-\beta\,\mathbb{D}_{\mathrm{KL}}(\pi_{\theta}\,\|\,\pi_{\mathrm{ref}})$
where now the advantage $A_{i}$ is computed as:
$A_{i}=\frac{R_{\mathrm{Multi\_Aspect}}(o_{i})-\mathrm{mean}\big(\{R_{\mathrm{Multi\_Aspect}}(o_{j})\}_{j=1}^{G}\big)}{\mathrm{std}\big(\{R_{\mathrm{Multi\_Aspect}}(o_{j})\}_{j=1}^{G}\big)+\varepsilon}$
The GRPO objective evaluates how well the current policy generates outputs compared to the previous policy. Probabilities under the current and previous policies are compared for importance sampling, while a clipping mechanism and KL-divergence penalty ensure stable updates and prevent excessive deviation from a reference policy. Each candidate’s advantage is computed relative to others in the same group using multi-aspect rewards, promoting the selection of higher-quality outputs without a separate value function. By integrating rewards for language consistency, text segmentation, object count, answer correctness, and output format, GRPO provides rich, fine-grained supervision for end-to-end multilingual multimodal reasoning.
Figure 3. Comparison between Qwen2.5-VL-7B and LaV-CoT. As illustrated, Qwen2.5-VL-7B demonstrates a step-by-step reasoning process; however, it fails to perform the reasoning in the target Arabic language and produces an incorrect final deduction. In contrast, LaV-CoT effectively follows its reasoning pipeline to produce the correct final answer.:
4. Experiment
In this section, we first provide an overview of the experimental framework, including implementation details, evaluation benchmarks, and the MLLMs used for comparison. We then present a comprehensive comparison of LaV-CoT with state-of-the-art approaches on multilingual benchmarks, followed by an evaluation across a variety of multimodal tasks. Finally, we conclude with ablation studies and qualitative visualization of multilingual cases, highlighting the exceptional reasoning capabilities of our model.
4.1. Experiment Setting
Implementation Details.
During the Supervised Fine-Tuning (SFT) stage, the model was trained for 1–3 epochs with a global batch size of 256 on 8 NVIDIA A100 (40GB) GPUs. The image resolution size is set to be 896 * 896. The learning rate was fixed at $2\times 10^{-4}$ and kept constant without decay, using the fused AdamW optimizer. Mixed-precision training with bfloat16 (bf16) and TensorFloat-32 (tf32) was enabled to improve computational efficiency while maintaining numerical stability. Gradient clipping with a maximum norm of 0.3 and a warm-up ratio of 0.03 were applied. Parameter-efficient fine-tuning was implemented via LoRA(Hu et al., 2021), targeting the transformer modules q_proj, k_proj, v_proj, o_proj, gate_proj, up_proj, and down_proj, with rank 32, alpha 64, and a dropout rate of 0.05. In the subsequent GRPO training stage, all reward components were uniformly scaled by 0.25 so that the total reward summed to 1, i.e.,
$R_{\mathrm{Multi\_Aspect}}=0.25\,R_{\mathrm{Lang}}+0.25\,R_{\mathrm{Count}}+0.25\,R_{\mathrm{Answer}}+0.25\,R_{\mathrm{Format}}$
The number of generations per iteration was set to 4, while other hyperparameters were kept consistent with the SFT stage.
Evaluation Benchmark.
For our experiments, MMMB(Sun et al., 2024), Multilingual MMBench(Sun et al., 2024), and MTVQA(Tang et al., 2024) serve as the primary evaluation benchmarks. Both MMMB and Multilingual MMBench focus on assessing the general multilingual visual question answering (VQA) capabilities of models, covering a wide range of question types and multimodal reasoning skills. While MTVQA is designed to specifically measure a model’s ability to align textual and visual information in text-centric VQA scenarios, with an emphasis on accurately recognizing and interpreting embedded text within images. Together, these datasets provide a complementary evaluation setting that captures both broad reasoning abilities and fine-grained multimodal text understanding. For evaluation, we adopt VLMEvalKit (Duan et al., 2024) with consistent configuration settings, thereby ensuring a fair and consistent comparison.
Comparison Models.
We adopt Qwen2.5-VL-3B as our backbone model and train it under the proposed LaV-CoT framework. For comprehensive evaluation, we compare against a diverse set of multilingual multimodal models (MLLMs), covering different model scales and accessibility levels. Specifically, we include: (1) open-source models of comparable size, such as Qwen2-VL-2B(Wang et al., 2024a), InternVL3-2B(Zhu et al., 2025), InternVL3.5-2B(Wang et al., 2025) and Qwen2.5-VL-3B (Bai et al., 2025) as the baseline; (2) larger open-source models with roughly double the parameters, including Monkey(Li et al., 2024b), LLaVA-OneVision(Li et al., 2024c), PARROT(Sun et al., 2024), DeepSeek-VL-7B(Lu et al., 2024), Qwen2-VL-7B(Wang et al., 2024a), Qwen2.5-VL-7B(Bai et al., 2025), InternVL3-8B(Zhu et al., 2025) and InternVL3.5-8B(Wang et al., 2025); (3) much larger proprietary frontier models, such as GPT-4o-0513(OpenAI, 2023) and Gemini-2.5-flash(DeepMind, 2023). In addition, we also report results of BlueLM-V-3B(Xiong et al., 2025), which, although closed-source, demonstrates strong multilingual and multimodal capabilities.
MMMB Multilingual MMBench MTVQA
Model Name en zh pt ar tr ru en zh pt ar tr ru ar de fr it ja ko ru th vi Overall
Qwen2-VL-2B 78.3 74.2 72.6 68.3 61.8 72.8 72.1 71.1 69.9 61.1 54.4 69.3 7.1 27.0 27.5 32.6 12.9 23.7 11.0 3.9 24.1 52.8
Qwen2.5-VL-3B 81.2 81.0 74.8 71.0 65.6 75.2 81.1 79.7 75.8 69.3 62.4 72.6 11.7 27.3 33.3 31.7 12.7 26.2 10.3 10.4 35.6 57.3
InternVL3-2B 81.9 78.3 75.4 68.6 62.9 74.6 81.3 77.8 75.9 66.4 59.5 70.7 5.5 34.0 41.1 39.7 19.6 31.5 9.4 16.5 28.5 57.4
InternVL3.5-2B 80.2 77.7 75.9 68.5 69.1 76.3 79.0 76.5 74.3 64.4 63.1 72.2 9.2 34.5 42.3 39.8 20.9 32.7 10.8 17.9 29.8 58.0
BlueLM-V-3B 17.3 39.5 44.7 32.2 23.5 34.0 9.2 20.3 22.9
Monkey 66.0 58.1 46.3 38.8 37.6 48.5 58.0 53.5 49.5 31.0 31.3 45.1 1.1 16.4 21.2 21.7 4.3 5.4 5.3 6.1 8.4 35.0
DeepSeek-VL-7B 72.6 65.9 64.4 49.7 49.0 67.5 70.7 64.0 62.6 48.0 47.9 65.5 1.4 18.4 20.4 18.0 5.1 7.0 1.6 3.5 7.2 43.9
GLM-4v-9B 69.2 62.8 61.5 47.2 46.9 64.3 67.9 61.3 60.0 46.1 45.7 63.0 7.0 31.4 39.3 37.9 11.1 13.4 8.1 8.2 26.6 46.2
LLaVA-OneVision 79.0 78.2 75.9 73.3 67.7 76.3 76.7 75.3 73.4 70.4 64.8 73.1 5.0 21.0 22.3 26.1 6.2 7.3 6.0 3.0 13.6 53.7
PARROT 80.1 80.0 79.6 76.5 75.0 79.9 78.0 77.1 76.7 75.9 74.0 77.7 14.3 30.1 34.6 36.7 16.3 20.2 9.5 9.9 27.8 59.7
Qwen2-VL-7B 83.9 82.4 81.2 79.0 74.7 82.4 81.8 81.6 79.1 75.6 74.5 79.3 16.2 30.5 35.8 36.9 16.8 30.1 11.1 11.3 28.7 61.6
Qwen2.5-VL-7B 85.0 83.6 82.1 83.3 76.4 83.2 85.3 85.8 83.0 80.2 75.7 82.9 17.8 30.5 33.3 37.2 18.5 38 13.6 15.2 44.8 64.4
InternVL3-8B 85.1 83.1 82.5 81.6 76.2 83.4 85.5 85.6 83.2 79.2 75.9 82.6 9.8 37.4 45.5 41.4 22.3 31.9 11.4 18.2 38.5 64.7
InternVL3.5-8B 84.9 83.0 81.4 79.6 77.4 83.1 84.5 85.7 80.9 82.8 75.8 82.3 16.1 38.2 44.1 40.6 22.8 33.7 11.6 20.2 39.8 64.9
GPT-4o-0513 84.9 84.3 82.8 82.3 79.0 83.3 87.6 88.2 85.5 85.6 82.9 86.2 21.3 35.1 42.2 37.2 19.9 35.1 15.9 26.0 39.6 66.6
Gemini-2.5-flash 85.2 84.5 83.1 81.8 79.6 84.0 88.3 89.0 86.1 84.2 85.8 88.1 21.6 36.8 41.7 43.0 19.8 36.3 15.8 26.9 41.7 67.4
LaV-CoT (SFT) 83.2 81.1 80.6 80.5 77.3 82.2 85.7 84.4 84.3 83.9 85.2 83.1 19.8 33.7 38.1 33.6 20.7 37.1 10.4 22.1 37.7 64.7
LaV-CoT (SFT + GRPO) 86.0 84.8 83.0 82.7 80.3 83.6 89.0 89.4 85.9 88.3 86.8 87.7 23.2 35.3 42.8 35.9 23.1 38.2 11.5 23.8 38.8 67.5
Table 1. Multilingual Vision-Language Reasoning Model Results. Results are presented for MMMB, Multilingual MMBench, and MTVQA across multiple languages.
5. Main Results
5.1. Quantitative Analysis
Table 1 presents the multilingual evaluation results of a wide range of vision-language models across three benchmarks: MMMB, Multilingual MMBench, and MTVQA. Several important observations can be drawn.
Overall and comparative performance. Large proprietary models such as GPT-4o and Gemini-2.5-flash achieve strong performance across most benchmarks, with overall accuracy of 66.6 and 67.4, respectively, establishing a strong upper bound for multilingual multimodal reasoning. Among open-source models, Qwen2.5-VL-7B, InternVL3.5-8B, and our proposed LaV-CoT variants deliver competitive performance, achieving overall scores of 64.4, 64.9 and 67.5, respectively. Notably, LaV-CoT (SFT + GRPO) surpasses all open-source baselines: on MMMB and Multilingual MMBench it performs strongly across most languages except for slightly lower scores on PT and RU, while on MTVQA it achieves particularly high accuracy for AR and KO.
Medium-sized models such as Monkey, DeepSeek-VL-7B, and GLM-4v-9B remain below 50 overall, highlighting the challenges of multilingual multimodal reasoning in low-resource settings. Smaller-scale models like Qwen2.5-VL-3B, InternVL3-2B and InternVL3.5-2B achieve modest improvements above 50, whereas larger open-source systems such as PARROT, LLaVA-OneVision, and Qwen2-VL-7B deliver substantially stronger results. Our LaV-CoT (SFT) model outperforms strong baselines including Qwen2.5-VL-7B and InternVL3.5-8B, and with reinforcement learning fine-tuning (SFT+GRPO), it achieves state-of-the-art performance, approaching or slightly exceeding proprietary models.
Impact of GRPO. Importantly, incorporating GRPO training further improves LaV-CoT. The LaV-CoT (SFT + GRPO) variant achieves an overall score of 67.5, outperforming all open-source models and even surpassing the proprietary Gemini-2.5-flash and GPT-4o on specific multilingual settings (e.g., Arabic, Turkish, Korean, etc.). This highlights the effectiveness of reinforcement learning with preference optimization in enhancing cross-lingual reasoning.
Benchmark-specific insights. On MMMB and Multilingual MMBench, LaV-CoT achieves robust improvements across multiple languages, particularly in Arabic and Turkish, where traditional baselines exhibit significant drops. On MTVQA, our model also delivers substantial gains in low-resource languages such as Arabic, Korean, and Vietnamese, showcasing stronger cross-lingual visual reasoning ability.
Overall, these results confirm that LaV-CoT, especially when enhanced with GRPO, achieves state-of-the-art performance among open-source multilingual multimodal models and narrows the gap with proprietary systems.
Figure 4. Training Reward Curves:
5.2. GRPO Training Reward Analysis
Figure 4 illustrates the smoothed reward curves during GRPO training, highlighting the evolution of four key reward components: Language Reward, Count Reward, Edit Distance Reward, and Format Reward. The Format Reward exhibits rapid initial improvement, ascending from approximately 0.125 to 0.25 within the first 850 steps before stabilizing, indicating the base model has desent instruction following capability and shows early convergence in format adherence. In contrast, the Language Reward experiences a brief decline from 0.13 to 0.11 by step 800, followed by a steady ascent to 0.25 around step 6500, with subsequent fluctuations reflecting ongoing refinement in linguistic quality. The Count Reward demonstrates a gradual rise from 0.06, including a plateau between steps 3000 and 4000, reaching stability at 0.24 post-step 6540, suggesting progressive optimization of quantitative accuracy. The Edit Distance Reward initially decreases from 0.13 to 0.10 by step 2600, then recovers to 0.22 by step 7500 and converges around 0.23, underscoring challenges in minimizing textual discrepancies early in training. Curves are smoothed using a moving average with a window size of 10, and shaded areas represent ±1 standard deviation, reflecting the variability in reward estimates. Overall, these dynamics demonstrate GRPO’s effectiveness in balancing multi-objective rewards, achieving high-fidelity convergence by the end of training.
Model $\text{OCR}_{avg}$ $\text{SR}_{avg}$ $\text{IR}_{avg}$ $\text{OL}_{avg}$ $\text{SIU}_{avg}$
Qwen2.5-VL-3B (Baseline) 83.7 68.5 74.4 72.2 75.0
Qwen2.5-VL-3B (Direct Train) 85.7 71.5 76.4 74.4 77.0
Qwen2.5-VL-3B (Text CoT) 87.7 73.5 78.4 76.1 79.0
LaV-CoT (SFT) 88.6 75.2 80.5 81.0 81.5
Table 2. Ablation study with different training methods.
Figure 5. Abliation Study on LaV-CoT GRPO Training.:
6. Abliataion Study
6.1. Impact of Language-Aware Visual CoT on Performance
Table 2 presents an ablation study comparing different training strategies for Qwen2.5-VL-3B. The baseline model is trained without CoT, while ”Direct Train” denotes standard supervised fine-tuning, and ”Text CoT” incorporates textual Chain-of-Thought reasoning. LaV-CoT (SFT) further integrates language-aware visual CoT, resulting in consistent improvements across all evaluation metrics. To evaluate the effectiveness of each method, we examine several representative tasks from Multilingual MMBench: Spatial Relationship (SR), Identity Reasoning (IR), Object Localization (OL), and Structured Image-Text Understanding (SIU). Notably, LaV-CoT achieves the highest scores across OCR, SR, IR, OL, and SIU, demonstrating the substantial benefits of incorporating structured multilingual visual reasoning into the model.
6.2. Rewards Ratio Ablation Study
To investigate the impact of different reward configurations on model performance, we conduct an ablation study by varying the reward ratios $(\alpha,\beta,\gamma,\delta)$ in the GRPO framework. Figure 5 illustrates the performance across nine representative languages on MTVQA. The baseline configuration $(0,0,0,0)$, which does not incorporate any reward signals, achieves the lowest performance, with an average accuracy of 28.90. Incorporating format-only rewards $(0,0,0,1)$ yields a minimal improvement over the baseline. Introducing partial structural rewards $(0,0,0.5,0.5)$ results in a moderate gain, increasing the average accuracy to 29.36. A more balanced reward setting $(0,0.33,0.33,0.33)$ further improves performance, achieving an average accuracy of 30.11. Finally, the fully balanced configuration $(0.25,0.25,0.25,0.25)$ delivers the best overall results, with an average accuracy of 31.13, demonstrating the effectiveness of jointly optimizing multiple reward dimensions. These findings highlight that carefully balancing the GRPO rewards is crucial for enhancing multilingual reasoning robustness.
7. Online A/B Test
To evaluate the real-world effectiveness of our proposed approach, we integrated LaV-CoT into a full-scale intelligent document processing system that provides multilingual visual question answering capabilities. The system operates in realistic enterprise environments, handling diverse document types, complex layouts, and multiple languages. Unlike traditional pipelines, which often rely on heuristic rules or single-stage reasoning models, our method employs structured multi-stage reasoning to sequentially process document content, summarize text segments, identify languages, generate object-level image captions, and perform step-by-step reasoning to answer complex queries.
We conducted a online A/B test comparing LaV-CoT with the existing production pipeline. Key business and user-centric metrics were tracked, including answer acceptance rate and user satisfaction scores collected through post-interaction feedback. The results demonstrate that LaV-CoT substantially outperforms the baseline: the answer acceptance rate increased by 8.7%, and the user satisfaction score improved by 12.4%, confirming that structured multi-stage reasoning enhances both accuracy and overall user experience in practical document understanding scenarios.
8. Conclusion
In this paper, we introduced LaV-CoT, a novel framework for multilingual multimodal question answering in real-world scenarios. By integrating language-aware visual Chain-of-Thought reasoning with structured supervision and a novel GRPO-based multi-aspect reward design, LaV-CoT effectively addresses the limitations of existing approaches that struggle with multilingual alignment, language inconsistency, and complex reasoning tasks. Our reward formulation jointly evaluates language consistency, structural accuracy, and semantic alignment, enabling more interpretable reasoning and significantly improving both model stability and generalization. Extensive experiments across diverse benchmarks demonstrate that LaV-CoT not only outperforms strong open-source baselines of similar scale, but also rivals much larger models and closed-source frontier systems. Furthermore, its deployment on a large-scale Intelligent Document Processing platform confirms its practical value, achieving notable improvements in accuracy, user satisfaction, and cost-effectiveness. Looking ahead, we plan to extend LaV-CoT to broader low-resource languages and domain-specific applications, while further exploring advanced reward modeling to enhance the robustness and inclusiveness of multilingual multimodal reasoning systems.
9. Limitations
Despite the promising results, LaV-CoT still has several limitations:
  • Sensitivity to multilingual input quality. The effectiveness of LaV-CoT is contingent on the quality of the multimodal inputs. When document images contain heavy language mixing across different scripts, the reasoning pipeline may struggle to maintain language consistency and semantic fidelity.
  • Coverage of low-resource languages. Our current training primarily relies on open-source multilingual VQA datasets that focus on high- and medium-resource languages. However, high-quality datasets for truly low-resource or less common languages remain scarce, and their collection and construction are necessary for broader inclusiveness.
  • Limited exploration of fast-slow reasoning integration. LaV-CoT is currently designed with a slow, multi-step reasoning pipeline to enhance interpretability. While effective, it does not yet incorporate fast thinking strategies or hybrid fast-slow reasoning mechanisms, which we plan to explore in future work to further improve efficiency and adaptability.
10. Ethical Use of Data
All examples in this work, including those in the Appendix, are drawn from open-source datasets or anonymized real-world data, ensuring no private information is disclosed.
Acknowledgements.
This work was supported by ZoloZ Pte. Ltd.
Appendix A Dataset Statistics
Table 3 provides an overview of the datasets used in our experiments. These benchmarks span multiple languages and cover a diverse range of tasks, ensuring a comprehensive evaluation of multilingual visual question answering.
Dataset Language Size
COCO2017 EN,ZH,PT,AR,TR,RU 73k
Visual-Genome EN,ZH,PT,AR,TR,RU 18k
GQA EN,ZH,PT,AR,TR,RU 15k
OCR-VQA EN,ZH,PT,AR,TR,RU 16k
TextVQA EN,ZH,PT,AR,TR,RU 4.7k
Llava-Pretrain EN,ZH,PT,AR,TR,RU 0.6k
MTVQA AR,DE,FR,IT,JA,KO,RU,TH,VI 21k
Table 3. The number of samples selected from each benchmark, along with the language categories covered.
Appendix B Prompt Design
B.1. Generator Prompt
Figure 6. Instructions for generating Vanilla CoT.:
Figure 7. Instructions for correcting error step.:
B.2. Evaluator Prompt
Figure 8. Instructions for evaluating cot step.:
B.3. Inference Prompt
Figure 9. Instructions for LaV-CoT inference on open-ended questions.:
Figure 10. Instructions for LaV-CoT inference on MCQ.:
Appendix C More Real-world Cases
Please see the next page for more cases.
Figure 11. Arabic Visa demo case.:
Figure 12. Indonesian ID demo case.: